OpenStudy (anonymous):

Cos_____=6/19 How do i set this up and solve it? I need to show my work

Nnesha (nnesha):

solve for what ?

OpenStudy (anonymous):

Find The missing value to the nearest hundredth

Nnesha (nnesha):

i guess they want s u to find an angle

Nnesha (nnesha):

invs cos (6/19)

OpenStudy (anonymous):

what does invs stand for?

Nnesha (nnesha):

\[\cos^{-1} \] arccos or invs

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so you just plug that into your calculator? If so how do i show that

Nnesha (nnesha):

hmm that's what i was thinking

Nnesha (nnesha):

hmm not a lot of work let's say x is missing value \[\cos x = \frac{ 6 }{ 19 }\] \[x = \cos^{-1} (\frac{ 6 }{ 19})\] and then use calculator

OpenStudy (anonymous):

oh ok i see now let me try

Nnesha (nnesha):

:-)

OpenStudy (anonymous):

71.59?

Nnesha (nnesha):

yeah..